OpenStudy (anonymous):

I need to find the volume of a cylinder, I am given the diameter (35') and the height (588')

OpenStudy (awy):

Volume formula for cylinder is pi × radius2 × height

OpenStudy (awy):

pi x radius ^ 2 x height its radius squared

OpenStudy (anonymous):

I tried that, I must be doing something wrong: v=pi x 12.5^2 x 35 v=pi x 156.25 x 35 v=490.87 x 35 v=17,180.59 meters I'm just not sure it's right

OpenStudy (anonymous):

oops feet not meters

OpenStudy (awy):

check your radius, it should be 35/2 to find the radius always.

OpenStudy (awy):

Answer should be 1924.23 ft.

OpenStudy (anonymous):

I wrote the height down wrong in my explination, height should be 588 Does this seem right? v=pi x r^2 x h v=pi x 17.5^2 x 588 v=pi x 306.25 x 588 v=962.11 x 588 v=565,722.30 ??? I'm so lost on this one...

OpenStudy (awy):

yup that's right; that's what I got too.